OpenStudy (anonymous):

If f(x) = −3x − 3 and g(x) = 2x + 1, find f/g(−2)

OpenStudy (jdoe0001):

I gather you meant to say f( g(-2) )

OpenStudy (anonymous):

The question said F over G and had (-2) next to it

OpenStudy (anonymous):

OK ! We have : \[\frac fg(x)=\frac{f(x)}{g(x)}\]

OpenStudy (jdoe0001):

$$ f(x) = −3x − 3\\ g(x) = 2x + 1\\ \cfrac{f(x)}{g(x)}\implies \cfrac{−3x − 3}{2x + 1} $$

OpenStudy (anonymous):

So : \[\frac fg(-2)=\frac{f(-2)}{g(-2)}=\frac{-3(-2)-3}{2(-2)-1}=\frac{6-3}{-4-1}=\frac{3}{-5}\]

OpenStudy (anonymous):

is the answer -1?

OpenStudy (anonymous):

@RedOrangeBlue Look at my replies above ! O.o